OpenStudy (anonymous):

An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. 240 mph 220 mph 200 mph 180 mph

OpenStudy (whpalmer4):

\[d = r t\]distance is rate * time speed of plane in still air is \(v\) speed of wind is \(v_w\) when the plane flies with the wind (tail wind), it goes \(v+v_w\) over the ground (still only \(v\) relative to the air, which is moving) when the plane flies into the wind (head wind), it goes \(v - v_w\) over the ground the distance out and back is the same, so we don't actually need to know it! \[d = (v+v_w)(2.5) = (v-v_w)(3.5)\] \[(v+v_w)(2.5) = (v-v_w)(3.5)\]\[2.5 v + 2.5v_w = 3.5 v - 3.5v_w\]Solve for \(v\) and plug in the value of \(v_w\).

OpenStudy (whpalmer4):

to check your work (essential!), find out how far the plane flies at v-40 for 3.5 hours and compare with how far it flies at v+40 for 2.5 hours. They should be the same!

OpenStudy (whpalmer4):

what do you get for the answer?

OpenStudy (anonymous):

how am I suppose to plug it in? I don't know how to solve that for the answer

OpenStudy (whpalmer4):

Add subtract multiply divide to get v on one side of the equals sign...then you can substitute the value of \(v_w = 40\) and get the value of \(v\)

OpenStudy (whpalmer4):

\[2.5v+2.5v_w = 3.5v - 3.5v_w\]what do you get if you subtract 2.5v from both sides? \[2.5v - 2.5v + 2.5v_w = 3.5v -2.5v - 3.5v_w\]\[2.5v_w=1.0v-3.5v_w\]Now can you solve that for \(v\)?

OpenStudy (anonymous):

6.0vw=1.0v

OpenStudy (whpalmer4):

yes, so now replace \(v_w\) with 40 and you have your answer...

OpenStudy (whpalmer4):

I don't like to replace numbers until all the algebra is done

OpenStudy (anonymous):

so 240

OpenStudy (whpalmer4):

yep! try it out as suggested to make sure it works. 3.5 hours at 200, 2.5 hours at 280